# Cyclicity of Multipliers on the Unit Ball of $\mathbb{C}^{n}$: A Corona-Based Approach

### v2

@@ Abstract.
We study the cyclicity of multipliers in Dirichlet-type spaces $D_{\alpha}(\mathbb{B}_{n})$. Specifically, we show that a multiplier $f$ analytic on a neighborhood of the closed unit ball, whose zero set on the unit sphere is a compact, smooth, complex tangential submanifold of real dimension $m\leq n-1$, is cyclic in $D_{\alpha}(\mathbb{B}_{n})$ if and only if $\alpha\leq\frac{2n-m}{2}$, where $m$ is the real dimension of the zero set of $f$ on the boundary. Our approach combines classical results on peak sets in $A^{\infty}(\mathbb{B}_{n})$ due to Chaumat and Chollet with a Corona-type theorem with two generators for the multiplier algebra.
@@ 1. Introduction
Given a Banach space $\mathcal{A}$ of holomorphic functions in a domain in $\mathbb{C}^{n}$, suppose that polynomials are dense in $\mathcal{A}$ and act as multipliers - that is, for every polynomial $p$, the product $ph\in\mathcal{A}$ whenever $h\in\mathcal{A}$. A function $f\in\mathcal{A}$ is said to be cyclic if the linear span of all polynomial multiples of $f$ is dense in $\mathcal{A}$. Equivalently, $f$ is cyclic if there exists a sequence of polynomials $p_{n}$ such that $\|fp_{n}-1\|_{\mathcal{A}}\to 0$ as $n\to\infty$. The problem of characterizing cyclic functions is highly sensitive to the choice of space and remains nontrivial, even in classical Hilbert spaces of holomorphic functions on the unit disk. Complete classifications, such as Beurling’s theorem for Hardy space [5], are exceptional. The Brown–Shields conjecture, proposed in 1984 for the Dirichlet space, remains unresolved [2]. Recent progress in cyclicity in spaces of holomorphic functions on the unit disk, particularly Dirichlet-type spaces, can be found in [3], [9], [11], and [7]. A comprehensive textbook treatment of Dirichlet-type spaces is also available in [8].
In several variables, a complete characterization of cyclic functions in Dirichlet-type spaces remains elusive even for classical domains such as the unit ball and the polydisk. However, significant progress has been made in identifying cyclic polynomials, with explicit characterizations obtained in the unit ball of $\mathbb{C}^{2}$ and the bidisk [4], [13], [12]. Here, we prove a cyclicity result for multipliers whose zero set on the boundary has a specific structure. Our main result is the following:
@@
Let $f\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))\cap\mathcal{O}(\overline{\mathbb{B}}_{n})$ be nonvanishing in the unit ball of $\mathbb{C}^{n}$, and suppose that $\mathcal{Z}(f)\cap\overline{\mathbb{B}}_{n}=M$, where $M$ is a compact complex tangential smooth manifold. Let $m\coloneqq\dim_{\mathbb{R}}(\mathcal{Z}(f)\cap\overline{\mathbb{B}}_{n})$, if $\alpha\leq\frac{2n-m}{2}$, then $f$ is cyclic in $D_{\alpha}(\mathbb{B}_{n})$.
A combination of the results in [17], which give a sufficient condition for non-cyclicity, and Theorem 1 leads to the following corollary.
@@
Our main tool for approaching the problem is the following result. An analogous version in one variable was proved in [10] using Tolokonnikov’s Corona Theorem [16], and a slightly more general statement in several variables was recently obtained in [1]. We follow the same strategy here, using Theorem 8 to establish the result. This, together with the construction of a cyclic function $g$ in the multiplier space of $D_{\alpha}(\mathbb{B}_{n})$, denoted by $\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$, dominated by $f$, will help us prove Theorem 1.
@@
Let $f,g\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))\cap\mathcal{C}^{\infty}(\overline{\mathbb{B}}_{n})$ and assume that on $\mathbb{B}_{n}$ we have $|g|\leq|f|$ then there exists a natural number $N$ such that:
@@
The organization of the paper is as follows. In Section 2, we develop the necessary tools to prove the main result. In particular, Subsection 2.1 introduces the notation and basic definitions; Subsection 2.2 is devoted to the proof of Theorem 3; and Subsection 2.3 presents several classical results of Chaumat and Chollet concerning peak sets of $A^{\infty}(\Omega)$, where $\Omega$ is a strictly pseudoconvex domain, and concludes with Corollary 10. These results are essential for constructing a cyclic function in the multiplier algebra with the desired zero set. In Section 3, we provide the proof of our main result, Theorem 1.
@@ 2.1. Basics and Notations
$\|f\|_{D_{\alpha}(\mathbb{B}_{n})}^{2}:=\sum_{|k|=0}^{\infty}(n+|k|)^{\alpha}\cdot\frac{(n-1)!}{k!\,(n-1+|k|)!}\,|a_{k}|^{2}<\infty,$
@@
The space $\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$ is a Banach algebra when equipped with the operator norm induced by multiplication. In particular, when $\alpha=0$, so that $D_{0}(\mathbb{B}_{n})=H^{2}(\mathbb{B}_{n})$, we have
@@
The following Lemma is a direct consequence of the fact that $D_{\alpha}(\mathbb{B}_{n})=A^{2}_{-(\alpha+1)}(\mathbb{B}_{n})$ in the sense of the generalization given to the definition of Weighted Bergman Spaces in [18].
Lemma 4.
(2) $f||^{2}_{\alpha}\simeq\int_{\mathbb{B}_{n}}(1-|z|^{2})^{-(\alpha+1)}|f(z)|^{2}\,dv(z)$
@@
Let $\alpha\in\mathbb{R}$, and let $k\in\mathbb{N}$ be such that $\alpha-2k<0$. If $f\in D_{\alpha}(\mathbb{B}_{n})$ and $R^{k}f\in H^{\infty}(\mathbb{B}_{n})$, then $f\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$.
@@
A function $f\in D_{\alpha}(\mathbb{B}_{n})$ is called a shift-cyclic vector if:
$f]:=\overline{\operatorname{span}\{pf:p\in\mathbb{C}[z_{1},\ldots,z_{n}]\}}=D_{\alpha}(\mathbb{B}_{n}).$
@@ 2.2. Proof of Theorem 3
In order to prove Theorem 8 we need the following classical result on the solution of the $\bar{\partial}$-problem by Lieb and Range [15].
@@
Let $\Omega\subset\subset\mathbb{C}^{n}$ be a strictly pseudoconvex domain with smooth boundary of class $\mathcal{C}^{k+2}$. Then for each $q=1,2,\ldots,n$, there exist linear operators
@@
    If $f\in\mathcal{C}_{(0,q)}^{k}(\overline{\Omega})$ with $\bar{\partial}f=0$, then for all $k=0,1,2,\ldots$,
    $\|T_{q}^{*}f\|_{\mathcal{C}^{k+1/2}(\overline{\Omega})}\leq C_{k}\|f\|_{\mathcal{C}^{k}(\overline{\Omega})}.$
The constants $C_{k}$ depend only on $k$ and $\Omega$, and are invariant under small $\mathcal{C}^{k+2}$-perturbations of $\Omega$.
Here, $\|\cdot\|_{\mathcal{C}^{k+1/2}(\overline{\Omega})}$ and $\|\cdot\|_{\mathcal{C}^{k}(\overline{\Omega})}$ denote the Hölder norms, where by $\|f\|_{k}$ and $\|f\|_{k+\varepsilon}$ we mean:
$\|f\|_{k}=\sum_{|\alpha|\leq k}\sup_{x\in\overline{\Omega}}|D^{\alpha}f(x)|,$
$\|f\|_{k+\varepsilon}=\|f\|_{k}+\sum_{|\alpha|=k}\sup_{x\neq y}\frac{|D^{\alpha}f(x)-D^{\alpha}f(y)|}{|x-y|^{\varepsilon}}.$
@@
Let $f_{2},f_{1}\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))\cap\mathcal{C}^{\infty}(\overline{\mathbb{B}}_{n})$ and there exists $\delta>0$ such that $|f_{1}|^{2}+|f_{2}|^{2}\geq\delta$ on $\mathbb{B}_{n}$ then there exists $g_{1},g_{2}\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))\cap\mathcal{C}^{\infty}(\overline{\mathbb{B}}_{n})$ and natural number $k$ satisfying:
@@
    $\|g_{i}\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\leq\frac{C}{\delta^{k}}$
for some constant $C>0$.
@@
Moreover, one easily observes that $\omega$ is a closed $(0,1)$-form, smooth up to the boundary, and all of its derivatives have $L^{\infty}$ norms bounded by constants of the form $\frac{C}{\delta^{k}}$, where $k$ depends on the order of the derivative. By Theorem 7, there exists a solution $\phi$ to the problem $\bar{\partial}\phi=\omega$, which is also smooth up to the boundary, and all of its derivatives have $L^{\infty}$ norms bounded by constants of the form $\frac{C}{\delta^{k}}$, again with $k$ depending on the differentiation order.
Now, we claim that the choice
$g_{1}=g_{1}^{0}-f_{2}\phi,\quad g_{2}=g_{2}^{0}+f_{1}\phi$
is the desired solution. Clearly, it satisfies (i). Moreover, by Lemma 4, a sufficient condition for $g_{1}$ and $g_{2}$ to be in the multiplier space of $D_{\alpha}(\mathbb{B}_{n})$ is that their radial derivatives up to a certain order are $L^{\infty}$-bounded. Hence, the norm estimate in (ii) follows. ∎
@@
    $\|g_{i}\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\lesssim\frac{1}{\delta(\lambda)^{k}}\asymp|\lambda|^{2k}$ for $i=1,2$.
Now define the standard quotient map
@@
$\pi(1-\lambda g)^{-1}=\pi(g_{1}).$
Thus, $\pi((1-\lambda g))^{-1}$, which is a holomorphic Banach-valued function of $\lambda$, is defined for every $\lambda$, so it is an entire function. Hence, we may write:
(3) $\|\pi((1-\lambda g))^{-1}\|=\left\|\sum_{j\geq 0}\lambda^{j}\pi(g^{j})\right\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))/[f]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}}=\|\pi(g_{1})\|\leq\|g_{1}\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\lesssim|\lambda|^{2k}.$
Now, by Liouville’s Theorem, we conclude that $\pi(g^{2k+1})=0$. In other words,
$g^{2k+1}]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\subseteq[f]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))},$

### v3

# arXiv:2508.05189v3[math.CV] 28 Feb 2026

@@ Abstract.
We study the cyclicity of multipliers in Dirichlet-type spaces $D_{\alpha}(\mathbb{B}_{n})$. Specifically, we show that a multiplier $f$ analytic on a neighborhood of $\overline{\mathbb{B}}_{n}$, whose zero set on the unit sphere is a compact, smooth, complex tangential submanifold of real dimension $m\leq n-1$, is cyclic in $D_{\alpha}(\mathbb{B}_{n})$ if and only if $\alpha\leq\frac{2n-m}{2}$. Our approach combines classical results on peak sets in $A^{\infty}(\mathbb{B}_{n})$ due to Chaumat and Chollet with a Corona-type theorem with two generators for the multiplier algebra.
@@ 1. Introduction
Given a Banach space $\mathcal{A}$ of holomorphic functions in a domain in $\mathbb{C}^{n}$, suppose that polynomials are dense in $\mathcal{A}$ and act as multipliers - that is, for every polynomial $p$, the product $ph\in\mathcal{A}$ whenever $h\in\mathcal{A}$. A function $f\in\mathcal{A}$ is said to be cyclic if the linear span of all polynomial multiples of $f$ is dense in $\mathcal{A}$. Equivalently, $f$ is cyclic if there exists a sequence of polynomials $p_{n}$ such that $\|fp_{n}-1\|_{\mathcal{A}}\to 0$ as $n\to\infty$. An immediate consequence of this definition is that, on a fixed domain, cyclicity becomes harder to achieve as the norm increases. The problem of characterizing cyclic functions is highly dependent on the choice of space and remains nontrivial, even in classical Hilbert spaces of holomorphic functions on the unit disk. Complete classifications, such as Beurling’s theorem for the Hardy space [5], are exceptional. The Brown–Shields conjecture, proposed in 1984 for the Dirichlet space, remains unresolved [2]. Recent progress in cyclicity in spaces of holomorphic functions on the unit disk, particularly Dirichlet-type spaces, can be found in [3], [9], [11], and [7]. A comprehensive textbook treatment of Dirichlet-type spaces is also available in [8].
In several variables, a complete characterization of cyclic functions in Dirichlet-type spaces remains elusive, even for classical domains such as the unit ball and the polydisk. However, significant progress has been made in identifying cyclic polynomials, with explicit characterizations obtained in the unit ball of $\mathbb{C}^{2}$ and the bidisk [4], [14], [12].
Here, we study Dirichlet-type spaces of the unit ball $D_{\alpha}(\mathbb{B}_{n})$, $\alpha\in\mathbb{R}$ , endowed with the norm $\|\cdot\|_{\alpha}$ (see Definition 7). Given $f\in\mathcal{O}(\overline{\mathbb{B}}_{n})$, a holomorphic functions defined on a neighborhood of the closed unit ball, we want to know for which $\alpha$ the function $f$ is cyclic in $D_{\alpha}(\mathbb{B}_{n})$. Note that when $\alpha\leq\beta$, $\|g\|_{\alpha}\leq\|g\|_{\beta}$, so as remarked above, if $f$ is cyclic in $D_{\beta}(\mathbb{B}_{n})$, it is cyclic in $D_{\alpha}(\mathbb{B}_{n})$ too.
If $f\in\mathcal{O}(\overline{\mathbb{B}}_{n})$ is non-vanishing in $\mathbb{B}_{n}$, it is known by [19, Theorem 1] that its zero set on the boundary satisfies
$\mathcal{Z}(f)\cap\partial\mathbb{B}_{n}=\bigcup_{i=1}^{k}M_{i},$
where the $M_{i}$ are smooth complex tangential manifolds (see Definition 19). If $\mathcal{Z}(f)\cap\partial\mathbb{B}_{n}$ is non-singular, then each $M_{i}$ is compact. The following theorem constitutes our main result and resolves the cyclicity problem for $f$ in this setting:
@@
Let $f\in\mathcal{O}(\overline{\mathbb{B}}_{n})$ be nonvanishing in $\mathbb{B}_{n}$, and suppose that $\mathcal{Z}(f)\cap\overline{\mathbb{B}}_{n}=M$, where $M$ is a compact complex tangential smooth manifold. Let $m\coloneqq\dim_{\mathbb{R}}M$, if $\alpha\leq\frac{2n-m}{2}$, then $f$ is cyclic in $D_{\alpha}(\mathbb{B}_{n})$.
A combination of Theorem 3 in [19], which gives a sufficient condition for non-cyclicity, and Theorem 1 leads to the following corollary.
@@
A consequence of Theorem 1.5 in a recent preprint by Knese, Pascoe, and Sola [13] is that the zero set in the sphere of an irreducible polynomial, non vanishing on $\mathbb{B}_{2}$, is a union of finitely many smooth compact complex tangential manifolds. Therefore, our main result provides an alternative way to characterize cyclic polynomials in $\mathbb{B}_{2}$, previously obtained by Kosiński and Vavitsas [14]. This follows from the fact that the product of two cyclic multipliers is cyclic (see page 13 of [18] for a proof of this fact).
The proof of our main result relies on two main ingredients. The first is a result of Chaumat and Chollet (Theorem 22) on peak sets for the smooth ball algebra
$A^{\infty}(\mathbb{B}_{n}):=\mathcal{O}(\mathbb{B}_{n})\cap\mathcal{C}^{\infty}(\overline{\mathbb{B}}_{n})$
(see Definition 21). This result allows us to construct a holomorphic function with the desired zero set together with sharp estimates to determine cyclicity. The second tool is the following result. An analogous version in one variable was proved in [10] using Tolokonnikov’s Corona Theorem [17]. We follow the same strategy, using Theorem 17 to obtain the result.
@@
Let $f,g\in A^{\infty}(\mathbb{B}_{n})$ and assume that on $\mathbb{B}_{n}$ we have $|g|\leq|f|$ then there exists a natural number $N$ such that:
@@
After completion of this work, Stefan Richter kindly informed us that a more general statement of Theorem 3 was already obtained in [1, Theorem 3.2] using a different method.
The organization of the paper is as follows. In Section 2, we develop the necessary tools to prove the main result. In particular, Subsection 2.1 introduces the notation and basic definitions; Subsection 2.2 is devoted to the proof of Theorem 3; and Subsection 2.3 presents a result of Chaumat and Chollet concerning peak sets of $A^{\infty}(\mathbb{B}_{n})$. In Section 3, we provide the proof of our main result, Theorem 1.
@@ 2.1. Basics and Notations
Definition 6.
Let $z_{0}\in\mathbb{C}^{n}$ and $E\subset\mathbb{C}^{n}$. We define
$\mathrm{dist}(z_{0},E)=\inf_{w\in E}|z_{0}-w|.$
@@
(2) $\|f\|_{\alpha}^{2}:=\sum_{|k|=0}^{\infty}(n+|k|)^{\alpha}\cdot\frac{(n-1)!\,k!}{(n-1+|k|)!}\,|a_{k}|^{2}<\infty,$
@@
In particular, when $\alpha=0$, $D_{0}(\mathbb{B}_{n})=H^{2}(\mathbb{B}_{n})$, the Hardy space. We will systematically use the characterization of Dirichlet-type spaces as generalized weighted Bergman spaces. Recall the classical definition of weighted Bergman spaces:
Definition 8.
Let $dv$ denote the $2n$-dimensional Lebesgue measure on $\mathbb{C}^{n}$. A function $f\in\mathcal{O}(\mathbb{B}_{n})$ belongs to the weighted Bergman space $A^{2}_{\beta}(\mathbb{B}_{n})$, where $\beta>-1$, if
$\|f\|^{2}_{A^{2}_{\beta}(\mathbb{B}_{n})}=\int_{\mathbb{B}_{n}}(1-|z|^{2})^{\beta}|f(z)|^{2}\,dv(z)<\infty.$
This definition was extended to all $\beta\in\mathbb{R}$ by Zhao and Zhu [20].
Definition 9.
  1. (1)
    Given $f\in\mathcal{O}(\mathbb{B}_{n})$, the radial derivative of $f$, denoted by $Rf$, is defined as follows:
    $Rf(z)=\sum_{i=1}^{n}z_{i}\frac{\partial f}{\partial z_{i}}(z).$
  2. (2)
    Given $\beta\in\mathbb{R}$ and $k\in\mathbb{Z}_{+}$ such that $\beta+2k>-1$,
    $f\in A^{2}_{\beta}(\mathbb{B}_{n})\mbox{ if and only if }\int_{\mathbb{B}_{n}}(1-|z|^{2})^{\beta+2k}|R^{k}f(z)|^{2}\,dv(z)<\infty.$
This definition is consistent with the classical one and does not depend on the choice of $k$. It is an important fact that $D_{\alpha}(\mathbb{B}_{n})=A^{2}_{-(\alpha+1)}(\mathbb{B}_{n})$. (Compare the power series norm in (2) with (20) in [20].) As a direct consequence we have:
Lemma 10.
(3) $\|f\|^{2}_{\alpha}\simeq\int_{\mathbb{B}_{n}}(1-|z|^{2})^{-(\alpha+1)}|f(z)|^{2}\,dv(z)$
@@
The space $\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$ is a Banach algebra when equipped with the operator norm induced by multiplication. In the well-known case of the Hardy space, we have
@@
As a corollary of Lemma 10, we have an easy sufficient condition to be a multiplier.
@@
Let $\alpha\in\mathbb{R}$, and let $k\in\mathbb{N}$ be such that $\alpha-2k<0$. If $f\in D_{\alpha}(\mathbb{B}_{n})$ and $R^{k}f\in H^{\infty}(\mathbb{B}_{n})$, then $f\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$. In particular, $\mathcal{O}(\overline{\mathbb{B}}_{n})\subset A^{\infty}(\mathbb{B}_{n})\subset\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$.
Proof.
Given $h\in\mathcal{O}(\mathbb{B}_{n})$ we have
$\left(\int_{\mathbb{B}_{n}}|h(z)|^{2}\,dv(z)\right)^{\frac{1}{2}}\leq|h(0)|+\left(\int_{\mathbb{B}_{n}}|Rh(z)|^{2}\,dv(z)\right)^{\frac{1}{2}}.$
Therefore, for $g\in D_{\alpha}(\mathbb{B}_{n})$ we may write
$\|fg\|_{\alpha}$ $\asymp|f(0)g(0)|+\left(\int_{\mathbb{B}_{n}}(1-|z|^{2})^{-(\alpha-2k+1)}\left|R^{k}(fg)(z)\right|^{2}\,dv(z)\right)^{\frac{1}{2}}$
$\leq C_{k}\sup_{0\leq l\leq k}\|R^{l}f\|_{L^{\infty}(\mathbb{B}_{n})}\left(|g(0)|+\sum_{l^{\prime}=0}^{k}\left(\int_{\mathbb{B}_{n}}(1-|z|^{2})^{-(\alpha-2k+1)}\left|R^{l^{\prime}}g(z)\right|^{2}\,dv(z)\right)^{\frac{1}{2}}\right)$
$\leq C^{\prime}_{k}\sup_{0\leq l\leq k}\|R^{l}f\|_{L^{\infty}(\mathbb{B}_{n})}\left(|g(0)|+\left(\int_{\mathbb{B}_{n}}(1-|z|^{2})^{-(\alpha-2k+1)}\left|R^{k}g(z)\right|^{2}\,dv(z)\right)^{\frac{1}{2}}\right)$
$\leq C^{\prime\prime}_{k}\sup_{0\leq l\leq k}\|R^{l}f\|_{L^{\infty}(\mathbb{B}_{n})}\|g\|_{\alpha}<\infty.$
So, $fg\in D_{\alpha}(\mathbb{B}_{n})$ and the proof is done. ∎
@@
The invariant (under the coordinate multiplication operators) subspace generated by $f$ is
$f]:=\overline{\operatorname{span}\{pf:p\in\mathbb{C}[z_{1},\ldots,z_{n}]\}}.$
A function $f\in D_{\alpha}(\mathbb{B}_{n})$ is called cyclic, or a cyclic vector, if $f]=D_{\alpha}(\mathbb{B}_{n}).$
@@ 2.2. Proof of Theorem 3
First, we need the following classical result on the solution of the $\bar{\partial}$-problem by Lieb and Range [16].
@@
Let $\Omega\subset\subset\mathbb{C}^{n}$ be a strictly pseudoconvex domain with $\mathcal{C}^{l+2}$ smooth boundary. Then for each $q=1,2,\ldots,n$, there exist linear operators
@@
    If $f\in\mathcal{C}_{(0,q)}^{l}(\overline{\Omega})$ with $\bar{\partial}f=0$, then for all $l=0,1,2,\ldots$,
    $\|T_{q}^{*}f\|_{\mathcal{C}^{l,1/2}(\overline{\Omega})}\leq C_{l}\|f\|_{\mathcal{C}^{l}(\overline{\Omega})}.$
The constants $C_{l}$ depend only on $l$ and $\Omega$.
Here, $\|\cdot\|_{\mathcal{C}^{l,1/2}(\overline{\Omega})}$ and $\|\cdot\|_{\mathcal{C}^{l}(\overline{\Omega})}$ denote the Hölder norms, obtained by:
$\|f\|_{{\mathcal{C}^{l}(\overline{\Omega})}}=\sum_{|\alpha|\leq l}\sup_{x\in\overline{\Omega}}|D^{\alpha}f(x)|,$
$\|f\|_{\mathcal{C}^{l,\varepsilon}(\overline{\Omega})}=\|f\|_{\mathcal{C}^{l}(\overline{\Omega})}+\sum_{|\alpha|=l}\ \sup_{x\neq y}\ \frac{|D^{\alpha}f(x)-D^{\alpha}f(y)|}{|x-y|^{\varepsilon}},\qquad 0<\varepsilon<1.$
Definition 16.
$A^{l,\varepsilon}(\mathbb{B}_{n})=\mathcal{O}(\mathbb{B}_{n})\cap\mathcal{C}^{l,\varepsilon}(\overline{\mathbb{B}}_{n})$.
@@
Let $f_{1},f_{2}\in A^{\infty}(\mathbb{B}_{n})$ and $\delta>0$ be such that $|f_{1}|^{2}+|f_{2}|^{2}\geq\delta$ on $\mathbb{B}_{n}$. Then, for any $\alpha\in\mathbb{R}$, there exist $g_{1},g_{2}\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$ and a natural number $k$ satisfying:
@@
    $\|g_{i}\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\leq\frac{C}{\delta^{k}}\|f_{1}\|_{\mathcal{C}^{k}(\overline{\mathbb{B}}_{n})}^{k}\|f_{2}\|_{\mathcal{C}^{k}(\overline{\mathbb{B}}_{n})}^{k}$.
for some constant $C>0$ dependent on $k$.
@@
Moreover, a direct computation shows that $\omega$ is a closed $(0,1)$-form and smooth up to the boundary.
Let $A=\bar{f}_{2}\,\bar{\partial}\bar{f}_{1}-\bar{f}_{1}\,\bar{\partial}\bar{f}_{2}$, $B=|f_{1}|^{2}+|f_{2}|^{2}$, and let $X_{1}X_{2}\ldots X_{l}$ be a differential operator of order $l$. For any $\{l_{1},\ldots,l_{m}\}\subseteq\{1,\ldots,l\}$, we have
$|X_{l_{1}}\ldots X_{l_{m}}A|\leq C_{m}\|f_{1}\|_{\mathcal{C}^{m+1}(\overline{\mathbb{B}}_{n})}\|f_{2}\|_{\mathcal{C}^{m+1}(\overline{\mathbb{B}}_{n})}\leq C_{m}\|f_{1}\|_{\mathcal{C}^{l+1}(\overline{\mathbb{B}}_{n})}\|f_{2}\|_{\mathcal{C}^{l+1}(\overline{\mathbb{B}}_{n})},$
and
$\left|X_{l_{1}}\ldots X_{l_{m}}\!\left(\frac{1}{B^{2}}\right)\right|\leq C^{\prime}_{m}\left|\frac{1}{B^{m+2}}\right|\|B\|_{\mathcal{C}^{m}(\overline{\mathbb{B}}_{n})}^{m}\leq C^{\prime\prime}_{m}\frac{1}{\delta^{m+2}}\|f_{1}\|_{\mathcal{C}^{m}(\overline{\mathbb{B}}_{n})}^{m}\|f_{2}\|_{\mathcal{C}^{m}(\overline{\mathbb{B}}_{n})}^{m}.$
This implies
$\|\omega\|_{{\mathcal{C}^{l}(\overline{\mathbb{B}}_{n})}}\lesssim\frac{1}{\delta^{l+2}}\|f_{1}\|_{\mathcal{C}^{l+1}(\overline{\mathbb{B}}_{n})}^{l+1}\|f_{2}\|_{\mathcal{C}^{l+1}(\overline{\mathbb{B}}_{n})}^{l+1}.$
By Theorem 15, there exists a solution $\phi_{l}$ to the problem $\bar{\partial}\phi_{l}=\omega$ such that
$\|\phi_{l}\|_{{\mathcal{C}^{l,\frac{1}{2}}(\overline{\mathbb{B}}_{n})}}\lesssim\frac{1}{\delta^{l+2}}\|f_{1}\|_{\mathcal{C}^{l+1}(\overline{\mathbb{B}}_{n})}^{l+1}\|f_{2}\|_{\mathcal{C}^{l+1}(\overline{\mathbb{B}}_{n})}^{l+1}.$
Now, let $k\in\mathbb{N}$ be such that $\alpha-2k<0$, we claim that the choice
$g_{1}=g_{1}^{0}-f_{2}\phi_{k},\quad g_{2}=g_{2}^{0}+f_{1}\phi_{k}$
is the desired solution. By construction, it satisfies (i). Moreover, by Lemma 10, a sufficient condition for $g_{1}$ and $g_{2}$ to be in the multiplier space of $D_{\alpha}(\mathbb{B}_{n})$ is that their radial derivatives up to order $k$ are $L^{\infty}$-bounded. Hence, the norm estimate in (ii) follows. ∎
@@
    $\|g_{i}\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\lesssim\frac{1}{\delta(\lambda)^{k}}\|1-\lambda g\|^{k}_{\mathcal{C}^{k}(\overline{\mathbb{B}}_{n})}\|f\|^{k}_{\mathcal{C}^{k}(\overline{\mathbb{B}}_{n})}\lesssim|\lambda|^{3k}$ for $i=1,2$.
Now let $f]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}$ be the closed ideal generated by $f$ in $\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$}. Define the standard quotient map
@@
$\pi\left(1-\lambda g\right)^{-1}=\pi(g_{1}).$
Thus, $\pi\left(1-\lambda g\right)^{-1}$, which is a holomorphic Banach-valued function of $\lambda$, is defined for every $\lambda$, so it is an entire function. Hence, we may write:
(4) $\Bigl\|\pi\left(1-\lambda g\right)^{-1}\Bigr\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))/[f]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}}$ $=\Bigl\|\pi(\sum_{j\geq 0}\lambda^{j}g^{j})\Bigr\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))/[f]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}}$
$=\Bigl\|\sum_{j\geq 0}\lambda^{j}\pi(g^{j})\Bigr\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))/[f]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}}$
$=\|\pi(g_{1})\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))/[f]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}}$
$\leq\|g_{1}\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\lesssim|\lambda|^{3k}.$
Now, by Liouville’s Theorem, we conclude that $\pi(g^{3k+1})=0$. In other words,
$g^{3k+1}]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\subseteq[f]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))},$